A 1.5 volt battery is connected to a 10 ohm resistor.

a. What current passes through the resistor?

b. What is the power developed by this resistor?



1
Expert's answer
2021-04-20T10:55:12-0400

"a. Current I =\frac {Voltage V} {Resistance R}"

"I= \frac{1.5} {10} = 0.15A"


b. "Power P = VI"

"P= 1.5\times 0.15= 0.225 W"